OpenStudy (anonymous):

find the complex zeros of x²+x+1.. I came up with -1+-3i/2?

OpenStudy (anonymous):

\[\frac{-1+\sqrt{3}}{2} and \frac{-1-\sqrt{3}}{2}\]

OpenStudy (anonymous):

Get it

OpenStudy (anonymous):

my book says -1/2 - √3/2 i, -1/2 + √3/2 are they the same answer just in a different form?

Directrix (directrix):

Your book answers are correct. They are not the same as the answers shown above because there is no i in those answers. Again, these are correct: -1/2 - √3/2 i, -1/2 + √3/2.